OpenStudy (anonymous):

Write an equation of a line in point-slope form with a slope of -3/4 and passing through the point (1, -5).

OpenStudy (texaschic101):

point slope form : y - y1 = m(x - x1) slope(m) = -3/4 (1,-5)....x1 = 1 and y1 = -5 now we sub y - y1 = m(x - x1) y - (-5) = -3/4(x - 1) --- simplify y + 5 = -3/4(x - 1)